OpenStudy (anonymous):

find the zeroes of the quadratic polynomial f(x)=abx^2+b^2x-acx+bc and verify the relationship between its zeroes and co-efficient

OpenStudy (anonymous):

plsssssssssssssssssssssssssssssssshelp

OpenStudy (anonymous):

abx^2+(b^2-ac)x-bc =abx^2+b^2x-acx-bc =bx(ax+b)-c(ax+b) =(ax+b)(bx-c) Therefore, zeroes are -b/a and c/b. Now the relationship can be verified.